OpenStudy (anonymous):

The variables x and y vary directly. Given x = 5 and y = 3, write an equation that relates the variables and find y when x = 2. A.y = 15x; y = 30 B.y = 5/3 x; y = 10/3 C.y= 0.6x; x = 1.2 D.y = 15/x; y = 7.5

OpenStudy (anonymous):

directly vary means y=kx now y=3 x=5 3/5=k equation y=(3/5)*x y=6/5 when x=2 C is the answer

OpenStudy (anonymous):

3/5=.6 y=.6x 5/6=1.2 1.2=x